OpenStudy (anonymous):

Solve by graphing: 3x-2y=6 x-y=2

OpenStudy (anonymous):

(Systems of equations xD)

OpenStudy (anonymous):

Do I have to change the 2 equations into slope intercept form?

OpenStudy (anonymous):

It would be best if you did

OpenStudy (anonymous):

okay, so then...

OpenStudy (anonymous):

solve for y in the 2nd equattion y = x-2 and plug into first one and solve for x after you get that plug the x value into y=x-2 to get the y value

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

yea thats right

OpenStudy (anonymous):

But then it's solve by graphing...so don't I have to graph to find the coordinates to solve it?

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

oh sorry you can graph both of them solve for y on both equations you should have 2 equation in this form y=mx+b b = intercept and m = slope the answer is when the 2 graphs intersect

OpenStudy (anonymous):

wait...but I got this random point smack in a middle of the box.

OpenStudy (anonymous):

So...how do I know what it is?

OpenStudy (anonymous):

hm.. thats odd it should intersect at (2,0) let me try it y= 3/2x -3 y=x-2